Old Country Links, Inc., produces sausages in three production departments—Mixing, Casing and Curing, and Packaging. In the Mixing Department, meats are prepared and ground and then mixed with spices. The spiced meat mixture is then transferred to the Casing and Curing Department, where the mixture is force-fed into casings and then hung and cured in climate-controlled smoking chambers. In the Packaging Department, the cured sausages are sorted, packed, and labeled. The company uses the weighted-average method in its process costing system. Data for September for the Casing and Curing Department follow:

Percent Completed
Units Mixing Materials Conversion
Work in process inventory, September 1 2 100 % 60 % 50 %
Work in process inventory, September 30 2 100 % 20 % 10 %


Mixing Materials Conversion
Work in process inventory, September 1 $ 4,624 $ 106 $ 628
Cost added during September $ 263,996 $ 21,028 $ 153,140


Mixing cost represents the costs of the spiced meat mixture transferred in from the Mixing Department. The spiced meat mixture is processed in the Casing and Curing Department in batches; each unit in the above table is a batch and one batch of spiced meat mixture produces a set amount of sausages that are passed on to the Packaging Department. During September, 119 batches (i.e., units) were completed and transferred to the Packaging Department.

Required:

1. Determine the Casing and Curing Department's equivalent units of production for mixing, materials, and conversion for the month of September.

2. Compute the Casing and Curing Department's cost per equivalent unit for mixing, materials, and conversion for the month of September.

3. Compute the Casing and Curing Department's cost of ending work in process inventory for mixing, materials, conversion, and in total for September.

4. Compute the Casing and Curing Department's cost of units transferred out to the Packaging Department for mixing, materials, conversion, and in total for September.

5. Prepare a cost reconciliation report for the Casing and Curing Department for September.
Answer #1

1.

Computation of equivalent units of production

Mixing

Materials

Conversion

Units transferred to the next department

119

119

119

Equivalent units in ending WIP inventory

          Mixing 2 units X 100% complete

2

          Materials 2 X 20%

0.4

          Conversion    2X 10%

0.2

121

119.4

119.2

2. Costs per equivalent units of production

Mixing

Materials

Conversion

Cost of beginning WIP inventory

4624

106

628

Cost added during the period

263996

21028

153140

Total costs (a)

268620

21134

153768

Equivalent units of production (b)

121

119.4

119.2

Cost per equivalent unit (a/b)

2220

177.0017

1290

3 & 4.

Costs of ending work in process inventory and units transferred out.

Mixing

Materials

Conversion

Total

Ending Work in process inventory

      Equivalent units

2

0.4

0.2

      Cost per equivalent unit

2220

177

1290

      Cost of ending work in process inventory

4440

70.8

258

4768.8

Units completed and transferred out

     Units transferred to next department

119

119

119

     Cost per equivalent unit

2220

177

1290

     Cost of units transferred out

264180

21063

153510

438753

5.

Cost reconciliation:

Cost of beginning WIP inventory = 4624+106+628 =5358

Cost added to production during the year = 263996+ 21028 + 153140 = 438,164